OpenStudy (anonymous):

Find the radius of a cone with a volume of 45.19 cm cubed and a height of 7.5 cm A)1.2 B)2.4 C)2.6 D)4.8

OpenStudy (anonymous):

Use the formula : \[Volume = (\pi r^2h)/3\]

OpenStudy (anonymous):

i have to find the radius not the volume

OpenStudy (anonymous):

you use @waterineyes formula and isolate the (r) algebraically

OpenStudy (anonymous):

Then, alter the formula like this: \[Radius = \sqrt{(3*Volume)/\pi h}\]

OpenStudy (anonymous):

i got 5.75669

OpenStudy (anonymous):

Take the square root of 5.75669... and get the answer...

OpenStudy (anonymous):

and then by squaring it.. my answer is B (:

OpenStudy (anonymous):

Yes you are damn true...